OpenStudy (anonymous):

What is the factored form of the following expression? x2 – x – 42 A. (x – 7)(x + 6) B. (x + 7)(x + 6) C. (x + 7)(x – 6) D. (x – 7)(x – 6)

OpenStudy (anonymous):



OpenStudy (anonymous):

x2 - x - 42 = x2 - 7x + 6x - 42 = x(x-7) + 6(x-7) = (x-7)(x+6)